OpenStudy (anonymous):

Integrate (sinx + cosx)^2?

OpenStudy (lgbasallote):

can you expand \[(\sin x + \cos x)^2\]

OpenStudy (lalaly):

\[(sinx+cosx)^2=\sin^2x+2sinxcosx+\cos^2x\]\[\int\limits{(sinx+cosx)^2}dx=\int\limits{(\sin^2x+\cos^2x)dx}+\int\limits{2sinxcosxdx}\]you know that\[\sin^2x+\cos^2x=1\]and\[2sinxcosx=\sin2x\]so the integral becomes\[\int\limits{(sinx+cosx)^2dx}=\int\limits{dx}+\int\limits{\sin(2x)dx}\]can u take it from here?

OpenStudy (anonymous):

Yes i think so thanks for the great explanation !

OpenStudy (lalaly):

:)

OpenStudy (anonymous):

I just didnt realize you had to expand :s but i get it now

OpenStudy (anonymous):

could someone explain to me how sin2x becomes sin^2x?

OpenStudy (lgbasallote):

@lalaly is so awesome isnt she? she's like SOOOO GREAT!!!

OpenStudy (lalaly):

no no no... sin(2x)=2sinxcosx

OpenStudy (lalaly):

lol ewgeeee thanks

OpenStudy (anonymous):

the answer in the text book says the answer is x + (sinx)^2 and i was confused how they got that

OpenStudy (lgbasallote):

x is the \(\int dx\) you know that right?

OpenStudy (anonymous):

yes i get that one but its the sin2x i dont get

OpenStudy (lgbasallote):

hehe that's all i got too :P but i do remember this =_=

OpenStudy (lgbasallote):

should be \[-\frac{\cos (2x)}{2}\] right?

OpenStudy (lalaly):

yeah

OpenStudy (anonymous):

:S it says x + (sinx)^2

OpenStudy (anonymous):

so i geuss typo or something?

OpenStudy (lalaly):

well wolfram approves my answer so i guess the answer u have is wrong

OpenStudy (anonymous):

okay thanks for all the help

OpenStudy (lalaly):

ur welcome

OpenStudy (lgbasallote):

well \[\int 2\sin x \cos x dx\] let u = \(\sin x\) \[\int 2udu = u^2 = \sin^2 x\]

OpenStudy (lgbasallote):

that's how your book did it

OpenStudy (anonymous):

okay i see thanks !

OpenStudy (lalaly):

lol nice@lgbasallote

OpenStudy (lgbasallote):

<tips hat>